OpenStudy (anonymous):

What is the vertex of the graph of y = −3x2 − 2x + 1?

OpenStudy (anonymous):

thanks

hero (hero):

@jpw, do you know how to find the vertex of the graph algebraically?

OpenStudy (anonymous):

no i need help

OpenStudy (anonymous):

no sorry

hero (hero):

The vertex of the quadratic function is \(\left(-\frac{b}{2a}\right), f\left(-\frac{b}{2a}\right)\)

OpenStudy (anonymous):

so how`would i plug the numbers in

hero (hero):

You can do that, or you can convert the equation to vertex form y = a(x - h) + k where (h , k) is the vertex.

hero (hero):

a = -3, b = -2 c = 1

OpenStudy (anonymous):

i like the first one better

OpenStudy (anonymous):

so it would be 2 divided by 9?

OpenStudy (anonymous):

@Hero

hero (hero):

How'd you get 9 ?

OpenStudy (anonymous):

2(-3) @Hero

hero (hero):

2(-3) = -6